From: NASA HQ
Posted: Friday, August 20, 2010
All ISS systems continue to function nominally, except those noted previously or below. Saturday – Crew off duty.
At wake-up, Fyodor Yurchikhin performed the regular daily early-morning check of the aerosol filters at the Russian Elektron O2 generator which Maxim Suraev had installed on 10/19 in gaps between the BZh Liquid Unit and the oxygen outlet pipe (filter FA-K) plus hydrogen outlet pipe (filter FA-V). [FE-5 again inspects the filters before bedtime tonight, currently a daily requirement per plan, with photographs to be taken if the filter packing is discolored.]
Upon wake-up, FE-2 Caldwell-Dyson finished her 4-day session of the medical protocol Pro K (Dietary Intake Can Predict and Protect against Changes in Bone Metabolism during Spaceflight and Recovery), her 4th onboard run, with the final controlled diet and diet logging after the urine pH spot test. Photo documentation was part of the activity. [Under Pro K, the crewmember measures and logs the pH value of a urine sample, to be collected the same time of day every day for 4 days. The crewmember also prepares a diet log and then annotates quantities of food packets consumed and supplements taken.]
For FE-6 Walker, it was the 3rd day of her 4-day Pro K controlled diet & diet logging session, after the urine pH spot test, her third onboard run.
Before her sleeptime, Shannon will set up the equipment for her next 24-hour urine collections of the Generic HRF (Human Research Facility) urine sampling protocol. [Based on crew feedback, new cold stowage hardware, and IPV (International Procedures Viewer) capabilities, the generic blood & urine procedures were created to allow an individual crewmember to select their payload complement and see specific requirements populated. Individual crewmembers will select their specific parameter in the procedures to reflect their science complement. Different crewmembers will have different required tubes and hardware configurations, so they should verify their choice selection before continuing with operations to ensure their specific instruction.] For Douglas Wheelock, it was the 2nd day of his 3rd session with the Pro K protocol after the urine pH spot test.
Caldwell-Dyson, Wheelock & Walker continued their current week-long activity with the post-wakeup experiment SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight), 3rd for Doug & Shannon, 7th for Tracy, transferring data from their Actiwatches to the HRF-1 (Human Research Facility 1) laptop. [To monitor his/her sleep/wake patterns and light exposure, the crewmember wears a special Actiwatch device which measures the light levels encountered by him/her as well as their patterns of sleep and activity throughout the Expedition, using the payload software for data logging and filling in questionnaire entries in the experiment’s laptop session file on the HRF-1 laptop. The log entries are done within 15 minutes of final awakening for seven consecutive days.]
Additionally, Shannon completed another session with the Reaction Self Test (Psychomotor Vigilance Self Test on the ISS) protocol. [The RST is performed twice daily (after wakeup & before bedtime) for 3 days prior to the sleep shift, the day(s) of the sleep shift and 5 days following a sleep shift. The experiment consists of a 5-minute reaction time task that allows crewmembers to monitor the daily effects of fatigue on performance while on ISS. The experiment provides objective feedback on neurobehavioral changes in attention, psychomotor speed, state stability, and impulsivity while on ISS missions, particularly as they relate to changes in circadian rhythms, sleep restrictions, and extended work shifts.]
The six station residents completed the regular weekly three-hour task of thorough cleaning of their home, including COL (Columbus Orbital Laboratory) and Kibo JPM (JEM Pressurized Module). ["Uborka", usually done on Saturdays, includes removal of food waste products, cleaning of compartments with vacuum cleaner, damp cleaning of the SM (Service Module) dining table, other frequently touched surfaces and surfaces where trash is collected, as well as the sleep stations with a standard cleaning solution; also, fan screens and grilles are cleaned to avoid temperature rises. Special cleaning is also done every 90 days on the HEPA (high-efficiency particulate air) bacteria filters in the Lab.]
FE-3 Kornienko completed the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]
Walker, Wheelock & Caldwell-Dyson filled out their weekly FFQs (Food Frequency Questionnaires) on the MEC (Medical Equipment Computer). [On the FFQs, NASA astronauts keep a personalized log of their nutritional intake over time on special MEC software. Recorded are the amounts consumed during the past week of such food items as beverages, cereals, grains, eggs, breads, snacks, sweets, fruit, beans, soup, vegetables, dairy, fish, meat, chicken, sauces & spreads, and vitamins. The FFQ is performed once a week to estimate nutrient intake from the previous week and to give recommendations to ground specialists that help maintain optimal crew health. Weekly estimation has been verified to be reliable enough that nutrients do not need to be tracked daily.]
At ~9:00am EDT, the crew conducted their regular WPC (Weekly Planning Conference) with the ground, discussing next week's "Look-Ahead Plan" (prepared jointly by MCC-H and TsUP timeline planners), via S-band/audio, reviewing upcoming activities and any concerns about future on-orbit events.
On TsUP Go, Misha Kornienko was to repressurize the ISS cabin atmosphere with N2 (nitrogen) from Progress M-06M/38P’s SrPK air supply tankage to make up total pressure.
CDR, FE-2, FE-3, FE-4 & FE-6 had their weekly PFCs (Private Family Conferences) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop), Tracy at ~6:25am, Alex at ~7:45am, Mikhail at ~8:15am, Shannon at ~2:00 pm, Doug at ~3:30pm EDT.
Jobs listed for Skvortsov & Kornienko today on the Russian discretionary “time permitting” task list were –
Another run of the GFI-8 "Uragan" (hurricane) earth-imaging program with the NIKON D2X digital camera photography with 800mm telelens (FE-3) for large glaciers of South Patagonia ice field (left of flight path) as targets, and
A session for Russia's EKON Environmental Safety Agency, making observations and taking KPT-3 aerial photography of environmental conditions on Earth using the NIKON D3X camera (CDR/FE-3).
The crew worked out on today’s 2-hr physical exercise protocol on the CEVIS cycle ergometer with vibration isolation (FE-2), TVIS treadmill with vibration isolation & stabilization (CDR, FE-3, FE-5), ARED advanced resistive exercise device (FE-2, FE-3, FE-4, FE-5, FE-6), T2/COLBERT advanced treadmill (FE-4, FE-6) and VELO ergometer bike with bungee cord load trainer (CDR). [T2 snubber arm inspection is no longer needed after every T2 session but must be done after the last T2 session of the day.]
CIR (Combustion Integrated Rack), MDCA/Flex: No report.
Commercial Photo (JAXA): No report. CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): No report.
CubeLab: Hardware installed.
CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Complete. DECLIC (Device for the Study of Critical Liquids & Crystallization, CNES/NASA): Currently DECLIC is OFF. The next run will start on 8/27 and run for 15 days. DomeGene (JAXA): Complete.
DOSIS (Active Measure, ESA): Impacted by ETCS failure, was inactive since 8/1 until reactivation on 8/17. Issue with missing DOSTEL1 files during last data downlink still under investigation.
EarthKAM (Earth Knowledge Acquired by Middle School Students): No report.
EDR (European Drawer Rack, ESA): No report.
ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.
EMCS (European Modular Cultivation System): No report.
ENose (Electronic Nose): No report.
EPM (European Physiology Module): Activated in support of CARD.
EPO (Educational Payload Operations, NASA):(Kids in Space): No report. EPO LES-2 (ESA): No report.
EPO COMMERCIAL (JAXA): Photo session was performed on 7/22. . EPO 3-min Video (JAXA): No report. EPO J-Astro Report (JAXA): No report. EPO Dewey’s Forest (JAXA): Closed out on 3/15. EPO Space Clothes (JAXA): Complete.
EPO Hiten (Dance, JAXA): No report.
EPO Moon Score (JAXA): No report.
EPO Try Zero-G (JAXA): No report.
EPO Kibo Kids Tour (JAXA): Complete. EPO Paper Craft (Origami, JAXA): No report. EPO Poem (JAXA): No report.
EPO Spiral Top (JAXA): No report.
ERB-2 (Erasmus Recording Binocular, ESA):[ERB-2 aims are to develop narrated video material for various PR & educational products & events, including a 3D interior station view.] No report.
Fish Scales (JAXA): Completed on FD7/ULF-4 and returned on STS-132.
FOAM STABILITY (ESA): No report.
FOCUS: No report. FSL (Fluid Science Laboratory, ESA): No report.
FWED (Flywheel Exercise Device, ESA): No report.
GENARA-A (Gravity Regulated Genes in Arabidopsis A/ESA): No report.
GEOFLOW: No report.
HAIR (JAXA): No report.
HDTV System (JAXA): Was delivered by HTV1.
Holter ECG (JAXA): No report.
HQPC (JAXA): Was delivered by 34P.
HREP (HICO/Hyperspectral Imager for the Coastal Ocean & RAIDS/Remote Atmospheric & Ionospheric Detection System/JAXA): HREP is operating nominally. The Ethernet link was restored so images stored onboard can be downlinked. Currently 1505 images have been taken to date.
ICE CRYSTAL (JAXA): Complete.
ICV (Integrated Cardiovascular): “Shannon, the ICV team is looking forward to next week's FD75 Exercise Echo session followed by your third "Robocop" ambulatory monitoring session later in the week.” IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.
INTEGRATED IMMUNE: No report.
InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): No report.
IRIS (Image Reversal in Space, CSA): No report.
IV Gen (Intravenous Fluids Generation): No report.
KID/KUBIK6: No report.
Kids in Micro-G: No report.
KUBIK 3 (ESA): No report.
LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): No report.
Matryoshka-2 (RSA): Acquiring science data.
Marangoni UVP (JAXA): 8th run was performed on 7/15-716. . MAXI (Monitor of All-sky X-ray Image, JAXA): Acquiring science data.
MDCA/Flex: See under CIR.
MEIS (Marangoni Experiment for ISS) in JAXA FPEF (Fluid Physics Experiment Facility): No report.
MISSE7 (Materials ISS Experiment): MISSE-7 is in a normal configuration except PEC B remains off due to thermal concerns. MISSE-7 is operating at reduced bandwidth until PEP R9 software patch is uploaded to the Payload MDM, NET 9/2/2010.
MPAC/SEED (JAXA): Completed on 19A FD4. MSG-SAME (Microgravity Science Glovebox): First carousel (6 samples) was successfully completed ahead of schedule. The carousel change-out is planned in the coming days. MSL (Materials Science Laboratory, ESA): No report.
MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC-1 “Pirs”.
MYCO 2 (JAXA): No report.
MyoLab (JAXA): Completed on 4/20. NANOSKELETON (Production of High Performance Nanomaterials in Microgravity, JAXA): Returned on 19A.
NEURORAD (JAXA): No report.
NEUROSPAT (ESA/Study of Spatial Cognition, Novelty Processing and Sensorimotor Integration): No report. NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete. NUTRITION w/REPOSITORY/ProK: No report. PADIAK: No report. PADLES (JAXA, Area PADLES 3/4; Passive Area Dosimeter for Lifescience Experiment in Space): No report.
PASSAGES (JAXA): No report.
PCDF-PU (Protein Crystallization Diagnostic Facility - Process Unit): No report.
PCG (JAXA, Protein Crystal Growth): No report.
PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): PCRF maintenance will be performed on 7/29-30. PMDIS (Perceptual Motor Deficits in Space): Complete.
POLCA/GRAVIGEN (ESA): Complete.
Pro K: No report.
RadGene & LOH (JAXA): Complete. RadSilk (JAXA): Samples were returned to ground on ULF3. RST/Reaction Self Test (Psychomotor Vigilance Self Test on the ISS): "Tracy, Wheels and Shannon, thank you for continuing to participate in Reaction Self Test! We have a data downlink scheduled next week. We look forward to receiving and reviewing your data. Tracy and Wheels, thank you again for taking time to complete Reaction Self Test around the EVAs. These data will be very helpful in assessing one of the prime specific aims for the experiment of evaluating the extent to which Reaction Self Test is sensitive to fatigue from work intensity. We greatly appreciate it!"
RYUTAI Rack (JAXA): PCRF checkout was performed on 7/29-30 (see PCRF). SAIBO Rack (JAXA): CBEF Rotor replacement was performed on 7/27-28.
SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): No report.
SCOF (Solution Crystallization Observation Facility, JAXA): No report. SEDA-AP (Space Environment Data Acquisition Equipment-Attached Payload, JAXA): Rebooting will be resumed on 8/20.
SHERE (Shear History Extensional Rheology Experiment): Complete. SLAMMD (Space Linear Acceleration Mass Measurement Device): No report.
SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Tracy, Shannon and Wheels: Thank you for completing your Sleep logs this week. Because the first Sleep log-entry activity did not make it onto the timeline for this past Monday, the last log-entry is scheduled for early next week.”
SMILES (JAXA): Survival Mode.
SODI/IVIDIL (Selectable Optical Diagnostics Instrument/Influence of Vibration on Diffusion in Liquids, ESA): No report.
SODI/DSC (Selectable Optical Diagnostics Instrument/Diffusion Soret Coefficient): No report.
SOLAR (Solar Monitoring Observatory): Sun observation window #31 started on 7/30. SOLAR was deactivated on 8/1 and was on survival power only due to the ETCS cooling loop shutdown. The Sun observation window #31 would have ended on 8/12. SOLAR could be put back on Feeder#1 on 8/18 in time for safing for thruster firing. Temperature of SOLACES could be monitored after SOLAR was back on feeder#1 and showed the value of -25.86 degC, so no switch ON of SOLACES was required. SOLAR was put back on Feeder#2 after SOLSPEC calibration. Currently out of Sun visibility window.
SOLO (Sodium Loading in Microgravity): No report. Space-DRUMS: No report. SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): No report.
SPICE (Smoke Point In Co-flow Experiment): No report.
SPINAL (Spinal Elongation): No report.
SWAB (Characterization of Microorganisms & Allergens in Spacecraft): No report.
TASTE IN SPACE (ESA): No report.
THERMOLAB (ESA): No report.
TRAC (Test of Reaction & Adaptation Capabilities): Planned.
VASCULAR (CSA): No report.
VCAM (Vehicle Cabin Atmosphere Module, NASA): No report.
VESSEL ID System (ESA): Deactivated on 8/1 due to the ETCS failure. Was reconnected to an alternative SUP outlet on 8/2. Acquiring data since. Command list file uplink performed on 8/18.
VESSEL IMAGING (ESA): No report.
VO2max (NASA): “Shannon, congratulations on your second VO2max with Thermolab session using the Portable PFS hardware. The protocol was successfully completed and the data were downlinked to the ground.”
VLE (Video Lessons ESA): No report.
WAICO #1/#2 (Waving and Coiling of Arabidopsis Roots at Different g-levels; ESA): No report.
YEAST B (ESA): No report.
CEO (Crew Earth Observations): "Through 8/17 the ground has received a total of 3,046 frames of E-24 CEO imagery for review and cataloging. “We are pleased to report your acquisition of the following targets with times corresponding to those of our daily CEO Target Request lists: Dead Sea, Israel – several fine sessions – requirements met for this target; Mauna Loa, Hawaii – a pair of excellent sessions in both context and detail – requirements met for this target; Brasilia, Brazil – again good imagery in both context and detail – requirements complete; Lisbon , Portugal – good mapping views – a single context view is still required; Lake Poopo, Bolivia – excellent detailed mapping view – only context views required now; and Brazilian Fires – views of the fires were acquired, but illumination and aspect of the fires were not suitable for publication. Thank you for your diligent efforts at CEO in the midst of all your EVA’s! Your handsome portrait of NASA’s JSC and the surrounding Clear Lake area was published on GSFC’s Earth Observatory website this past weekend. NASA centers are a favorite target of both Shuttle and ISS crews, but your detailed view acquired under cloudless skies on the afternoon of 7/31 is the best JSC photo that we have seen recently. Kudos to the crew!”
CEO (Crew Earth Observation) photo targets uplinked for today were South African fires(this is the burning season in tropical Africa. Angola is known as the country in which natural and set fires are the most prevalent in the world. Looking right of track, from near nadir to well off track to document fires and smoke palls. Fires and smoke should have been visible for ~4 mins from this start time. Combined with oblique look angles, present conditions of lower illumination can help to reveal smoke in the atmosphere. Including any recognizable features [such as lakes, and areas with less smoke] so that the locations of ISS CEO images can be determined. Astronaut images of fires and smoke have been used in published scientific studies of global biomass burning), Harare, Zimbabwe (looking just left of track for this capital city. Harare is relatively simple to locate. The long geological feature known as the Great Dyke is a prominent, dark linear that extends for hundreds of km. Reservoirs that supply Harare with water have been constructed against the hills of Great Dyke: these reservoirs are the next visual cues), and Santiago, Chile (the city lied immediately north of track and contrasts with the surrounding agricultural landscape. Prior astronaut images show significant expansion of the city southward towards the Rio Maipo, and elsewhere, in the 15 years between the images).